OpenStudy (anonymous):

solve for x Solve for x: 125^5x−9=(1/625)^9x−7

jhonyy9 (jhonyy9):

\[(5^{3})^{5}x-9=(5^{-4})^{9}x -7\]

jhonyy9 (jhonyy9):

\[5^{15}x -9 = 5^{-36}x -7 \]

jhonyy9 (jhonyy9):

\[5^{15}x - 5^{-36}x = 2\]

jhonyy9 (jhonyy9):

\[x(5^{15} -5^{-36})=2\]

jhonyy9 (jhonyy9):

\[x= 2/(5^{15} -5^{-36})\]